If fixed costs are $277,000, the unit selling price is $29, and the unit variable costs are $18, what is the break-even sales (units) if fixed costs are reduced by $43,700?

16,967 units

31,814 units

21,209 units

25,451 units

If fixed costs are $719,000 and variable costs are 63% of sales, what is the break-even point in sales dollars?

$1,943,243

$2,662,243

$1,171,970

$452,970

If fixed costs are $272,000, the unit selling price is $123, and the unit variable costs are $74, what is the break-even sales (units)?

5,551 units

2,211 units

1,381 units

3,676 units
Answer #1

Contribution margin=Sales-Variable costs

1.Contribution margin=(29-18)=$11

Breakeven=Fixed cost/Contribution margin

=(277000-43700)/11

=21209 units

2.Contribution margin ratio=(100-63)=37% of sales

Hence breakeven=fixed cost/Contribution margin ratio

=(719000/0.37)=1943243(Approx).

3.Contribution margin=(123-74)=$49

Hence breakeven=272000/49=5551 units(Approx).